# Predicting liquid properties and behavior via droplet pinch-off and machine learning

### v1

@@ Abstract
We employed unsupervised clustering on the feature representations extracted by the convolutional neural network to identify intrinsic groupings in the droplet dynamics, based on the droplet images in our validation dataset. From the behaviour of the Silhouette and DBI metrics across candidate values of $K$, we selected $K=5$ for both K-Means and Gaussian Mixture Models-GMM, corresponding to a maximum Silhouette Score. In the case of the DBI extending $K$ beyond this value did not result in large improvement, and, as previously noted, a smaller number of clusters is preferable for insight into the underlying physics.
@@
We computed the average decoded shape of all droplets within a cluster to visualize the typical droplet morphology for that cluster. Figure 4c shows these mean shapes for the GMM and K-Means partitions. Both methods readily and consistently identify Cluster 4 as a bulb attached to a very long filament, and Cluster 0 as a wide and rounded bulb coming from a-small-in-comparison nozzle (with no connecting filament). The other three clusters only differentiate in the relative size of the nozzle and its connecting filament. Cluster 1 consists of elongated droplets connected to a small nozzle through a short filament, Cluster 3 is a droplet connected, through a filament, to a equal-in-size nozzle, and Cluster 2 is a droplet connected to a intermediate-in-size filament and nozzle. The close agreement between the average shapes from the two methods confirms that both clustering approaches capture coherent and meaningful droplet shape features.
Refer to captiona)
Refer to captionb)
Refer to captionc)
Refer to captiond)
@@
Figure 5: Clustering of droplet breakup regimes in the BoBoOhOh and ReReOhOh spaces. Results for K-means is seen in (a) and (b), while GMM is found in (c) and (d).
@@
A central question arising from the clustering is whether the unsupervised method effectively identifies, or differentiates, the underlying regimes and the various other factors influencing the dripping and break-up dynamics. Following previous works (25), that have classified the liquid break-up dynamics in terms of scaling arguments, Figure 5a and Figure 5c shows the cluster data projected onto the log-log ($Re$, OhOh) plane for GMM and K-Means. Importantly, some clusters are readily physically interpretable. Indeed, decreasing $Oh$ indicates weakening viscous effects relative to inertial-capillary stresses, while increasing $Re$ reflects stronger inertia at fixed viscosity. Both methods clearly separate Cluster 4 and Cluster 0. Cluster 4 lies predominantly in the upper-left quadrant (high $Oh$, low $Re$), consistent with viscous-damped pinch-off with limited necking and slow recoil, and visually associated with very long filaments. In contrast, Cluster 0 concentrates in the lower-right region (low $Oh$, high $Re$), characteristic of inertial–capillary breakup producing small, nearly spherical droplets with weak viscous damping. Further analyses can be carried out by reflecting on the distribution of liquid formulations among the clusters, as summarized in the Supplementary Material. A clear pattern emerges from these analyses: high viscosity fluids, such as glycerin-rich solutions and high cSt silicone oils, are consistently grouped into Cluster 4 by both GMM and K-Means. These samples correspond to the most elongated droplets, supported by the high ratios $L_{2}/L_{1}$ and the mean shapes shown in Figure 4c. In contrast, low-viscosity and low-surface-tension fluids are more frequently assigned to Cluster 0. These groups correspond to rounder, more stable droplets with minimal tailing.
@@
The lack of a common denominator among data clusters 1, 2, and 3 suggests that a further factor, apart from the flow rate, the nozzle size, and the liquid characteristics, is playing a role in driving the dripping dynamics. Fortunately, dripping is well-known to be governed by a third dimensionless parameter, the Bond number (25). Figure 5a and 5c shows the data and clustering in the BoBoOhOh space, where all five clusters are well separated, indicating distinct regimes governed by the relative importance of gravity ($Bo$) and viscosity ($Oh$). Clusters 0-3 occupy a narrow range of $Oh$ but progressively shift toward larger $Bo$. This increase in $Bo$ indicates that these clusters primarily reflect the gradual strengthening of gravitational forces relative to capillarity, while viscous effects remain comparatively constant. Cluster 0 corresponds to the capillary-dominated regime, where droplets are nearly spherical and detach quasi-statically. The gravitational effect strengthens among Clusters 1-3; even though they have comparable viscosity, the increasing gravitation elongates the filaments. In contrast, Cluster 4 deviates markedly from this $Bo$-driven sequence; it spans to much higher $Oh$ values while maintaining moderate $Bo$, this corresponds to high viscous effects in which surface tension and gravity are both resisted by strong viscous damping.
@@
The clustering was performed on a feature vector that concatenates droplet-shape descriptors with the experimental settings (nozzle diameter and mean flow speed) and does not explicitly include fluid properties. The data separation in ReReOhOh and BoBoOhOh therefore suggests that these features implicitly encode the inertia–viscosity–gravity balance. Notably, the physical properties of the liquids were not input into the clustering. Thus the separation in Re-Oh and Bo-Oh space indicates that our method successfully uncovered the underlying physics linking the experimental settings to the observed droplets.The strong mapping between physical composition and cluster assignment provides further evidence that the unsupervised clusters are not only mathematically separable but also physically interpretable.
Conclusions
Previous studies have demonstrated that the simultaneous measurement of surface tension and viscosity is possible using specialized techniques (54), but most traditional characterization protocols rely on separate tensiometers and viscometers, which often demand substantial sample volumes (55) and/or are limited by its measurement range (16). Our approach provides a single-step measurement of both surface tension and viscosity, alleviating the need for multiple instruments (56). Our results demonstrate a generalizable route to rapid, multi-parameter characterization of liquids from a single pinch-off snapshot, which is suited for integration into automated or in-line monitoring systems. The findings open up new possibilities in inkjet printing, e.g. real-time characterisation of ink viscosity and/or surface tension during jetting (42, 57), or on biomedical diagnostics with rapid small-volume property checks (58), of on other industrial processes with on-the-fly materials characterization (59).
By making our dataset (and code) openly available to the community, we expect that the database and models will be extended and potentially improved by including numerically obtained results and with additional experiments under conditions not attainable in our facilities, thereby broadening validation and improving generalization.
Materials and Methods
Experimental setup
[…]
Liquid materials and characterisation
[…]
The physical properties, i.e. density, viscosity, and surface tension, of the test liquids were characterised at (working) room temperature of (21.0 $\pm$ 1.0 Celsius). Density was determined gravimetrically by weighing a known volume of liquid using an analytical balance. Viscosity was measured on a rotational rheometer (Anton Paar MCR 102e) equipped with a cone–plate geometry at a constant shear rate of $50\,\text{s}^{-1}$. Each sample was measured twenty times with the mean value taken as the working viscosity; the typical error in the viscosity measurement is $\delta\mu_{true}=\pm$ 0.10 mPa s. For liquids of low and moderate viscosity ( $\mu\lesssim 60\penalty 10000\ \mathrm{mPa\,s}$)), the surface tension was evaluated via the maximum bubble pressure method using a SITA pro line T15 tensiometer at bubble lifetimes from 1.0 s to 15.0 s. Within this range, surface tension was found to be independent of surface age, with a standard error of $\delta\sigma_{true}=\pm 0.2$ mN/m. As noted in the introduction, bubble-pressure tensiometers are known to provide unreliable readings for highly viscous liquids due to viscous and hydrodynamic effects(16, 60). Accordingly, for glycerin and water mixtures of a concentration $\geq 80$ wt% glycerin, and for 350 cSt and 1000 cSt silicone oils, surface tension values were taken from the manufacturers’ data-sheets. Liquid property data is presented in the Supplementary Material. We note that viscosity values span more than three orders of magnitude while surface tension triples in value. The domain of our data, parametrized in terms of $Re$ and $Oh$, can be found in Fig. 1b. As can be seen there, our conditions cover four orders of magnitude in the Ohnesorge number and five in Reynolds, and include simple dripping, and jetting as defined in previous works (33).
Image analysis
High-speed images were visually analysed to identify the nearest snapshot before break-up; this was done for each experiment. We carried out 840 conditions/experiments, each producing an image (snapshot) showing the break-up of the droplet. Each image was processed by custom MATLAB code to obtain the droplet profile and its contour. The code works by first subtracting a pre-recorded background (image without a droplet, just the background) from the droplet snapshot to enhance the interfacial contrast. The code then employs a global threshold function (Otsu’s method) to obtain a binary image(61). This is followed by steps to remove image artefacts and holes. The largest connected region in the binary image is then identified as the droplet, and its exterior boundary is extracted as the raw contour. Finally, the boundary is smoothed and resampled uniformly. The needle outer diameter was used as size reference. Using this calibration, lengths, such as the droplet diameter $D_{d}$ at pinch-off, could be calculated and recorded for each experiment for further analysis; further details can be found in the Supplementary Material.
Machine learning methodology
A common foundation across both approaches is the use of a Convolutional Neural Network Autoencoder (CNN-AE) (46) to extract low-dimensional representations of the droplet snapshots, in the form of a latent shape vector z. The primary motivation for employing a CNN is its ability to filter and de-noise raw image data while reducing the input dimension. These features are then used as model inputs in the supervised learning tasks, or as clustering coordinates in the unsupervised learning analyses. Our dataset is constructed from the droplet boundary-only images previously described. We note the experimental images were obtained as high-resolution monochromes, but they were neither square nor of the resolution required. Consequently, to ensure consistent input sizes for the autoencoder, we padded the original images with a white background to transform them into square images of a $875\times 875$ ($x\in\mathbb{R}^{875\times 875}$) size; this padding preserves the original resolution and aspect ratio of the droplet shapes without any downsampling or interpolation. CNNs require that all images within a batch share the same spatial dimensions; padding was therefore applied to obtain square inputs, standardizing image size and avoiding distortions caused by rescaling.
From our dataset of 840 samples we randomly selected 600 samples for training, 120 samples for validation, and 120 samples for the testing of the autoencoder and the other methods. At each step, these groups of samples remained isolated from each other to avoid data leakage. The network was trained using the Adam optimizer (62) with a learning rate of $2.5\times 10^{-4}$. An early stopping criterion was applied: if the validation loss did not decrease for 30 consecutive epochs, training was terminated. Under these conditions, the training converged after 140 epochs with a batch size of 32. Loss curves for the autoencoder are provided in Supplementary Material. It should be noted that the data used for validation are not the data used for final, out of sample testing; the test loss is only evaluated once training is completed and provides an unbiased measure of the final model performance. A learning rate scheduler reduces the learning rate by a factor of 0.5 after a quarter of the training epochs. To improve generalization, random Gaussian noise with a standard deviation of 0.02 is added to the input images during training. The model uses binary cross-entropy (BCE) loss, as the output is normalized. Mathematically, the autoencoder consists of an encoder $E_{\theta}$ and decoder $D_{\phi}$, parameterized by $\theta$ and $\phi$, respectively. Given an input image $x$, the BCE reconstruction loss is defined as
$\mathcal{L}_{\text{BCE}}(x,D_{\phi}(E_{\theta}(x)))$ $=-\frac{1}{N}\sum_{i=1}^{N}\Big[x_{i}\log(D_{\phi}(E_{\theta}(x))_{i})$ (1)
$\quad+(1-x_{i})\log(1-D_{\phi}(E_{\theta}(x))_{i})\Big],$
where $x$ is the ground truth image, $D_{\phi}(E_{\theta}(x))$ is the reconstructed image, both with pixel values in $[0,1]$, and $N$ is the total number of pixels. This loss encourages the model to accurately predict pixel intensities in the reconstructed image (63). The relative reconstruction error was computed as
$\text{RE}(x,D_{\phi}(E_{\theta}(x)))=\frac{\lVert x-D_{\phi}(E_{\theta}(x))\rVert_{2}}{\lVert x\rVert_{2}}.$ (2)
The validation set was then used to evaluate the reconstruction performance of the autoencoder; the relative reconstruction error as a function of the latent dimension can be found in the Supplementary Material. Our results indicate that the relative error decreases rapidly as the latent dimension $d$ (the length of vector z) increases, and plateaus around $d=14$, indicating that adding more dimensions does not improve accuracy. Consequently, in our methods, we set the autoencoder to produce a vector z of size $d=14$. We note that we have found a low-dimensional representation with a dimension in $\mathcal{O}(10^{1})$, starting from an input dimension of $\mathcal{O}(10^{5})$; thus a reduction of four orders of magnitude is achieved.
Predicting liquid properties & droplet shape with supervised learning
Four supervised models were used to infer either one or two fluid properties (viscosity and/or surface tension) from images, or to predict the droplet shape from a set of known physical parameters (model diagrams are found in the Supplementary Material). All supervised learning models were implemented using two types of algorithms: MLP and XGBoost (64), which differ in their learning paradigms. Neural networks learn smooth continuous mappings, while XGBoost captures nonlinear feature interactions through ensembles of decision trees—allowing a complementary comparison between deep and tree-based methods. Models 1–3 combine the autoencoder latent vector z with physical parameters (nozzle diameter, flow rate, density, viscosity, and surface tension) to predict fluid properties. Model 4 performs the inverse mapping, predicting droplet shape from the same physical parameters, thereby establishing a bidirectional link between geometry and material behaviour. All supervised learning models are implemented as fully connected feedforward neural networks. As for the case of the autoencoder, loss curves for the training of these models can be found in the Supplementary Material. In addition, the same train, validation, and test split as used for the autoencoder was used here, with the same strict separation of these sets, in order to prevent data leakage. Each model was trained using the Adam optimizer with an initial learning rate of $10^{-3}$ and the mean-squared-error (MSE) loss function,
$\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{y}_{i}-y_{i}\|_{2}^{2}.$ (3)
where $N$ denotes the number of samples in a batch. For Models 1–3, the output values were log-transformed during training to improve numerical stability. A batch size of 32 was employed, and a StepLR scheduler was applied with a step size of 100 epochs and a decay factor of $0.8$. Early stopping with a patience of 100 epochs was applied based on validation loss, and the model weights corresponding to the lowest validation loss were restored for testing. For targets with large dynamic range (e.g., viscosity), the outputs were log-transformed during training to improve numerical stability, and then mapped back to the physical scale for evaluation. Models 1 and 2 are formulated as single-task regressors. In Model 1, the inputs consist of the droplet shape latent vector obtained from the CNN encoder, together with the density, diameter, flow rate and the surface tension; the network is trained to predict the viscosity $\mu$. Model 2 receives the same latent vector, plus density, diameter, flow rate, and the viscosity $\mu$, to predict the surface tension $\sigma$. These two models served as baselines, inferring one unknown physical property of the example given the droplet shape (via the encoder representation) and the known other physical properties. Model 3 extends this approach to a multi-task setting. Here, the inputs are only the latent shape vector and the control parameters of density, nozzle size and flow rate, and the network simultaneously outputs both viscosity $\mu$ and surface tension $\sigma$.This design allows the model to learn a shared network structure for the two correlated fluid properties, and provides a more realistic scenario in which neither property is known a priori. Together, Models 1 to 3 provide a progressive comparison between single-task and multi-task learning for fluid property prediction. The performance of the models was assessed by the coefficient of determination $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (4)
where $y_{i}$ denotes the true (experimentally obtained) values, $\hat{y}_{i}$ the (ML-obtained) predicted values, and $\bar{y}$ the mean of the true values. A detailed summary of model architectures, input compositions, and output targets is provided in the Supplementary Material.
Uncovering natural groupings in droplet behaviour with unsupervised learning
In addition, we have employed unsupervised learning to directly uncover naturally-occurring group behaviour in droplet dynamics from the observed shapes and imposed flow conditions, without requiring prior knowledge of the underlying fluid properties.
This work is motivated by experimental scenarios where viscosity or surface tension are not readily measurable, yet droplet morphology and control parameters are accessible, e.g. in systems where liquid volumes are too small. Specifically, we extracted low-dimensional latent features $z\in\mathbb{R}^{14}$ from the encoder, with the latent dimension $d=14$ adopted from Section 3.1, where it was identified as the optimal representation size. These latent features were combined with two directly controlled experimental parameters, nozzle inner diameter and flow rate, for clustering. Notably, we deliberately excluded any information related to the fluid properties, as these quantities are often difficult to measure in practice. We applied both K-Means and GMM with cluster numbers $K$. To identify optimal settings, we systematically tested $K\in\{2,7\}$. The choice of the upper bound $K=7$ reflects a practical consideration: while larger values of $K$ might further partition the data, excessively fine-grained clustering may produce groups which lack clear physical interpretation in the context of droplet dynamics. Thus, restricting $K$ to moderate values ensures a balance between capturing heterogeneity in the dataset and maintaining physically meaningful cluster assignments.
In order to assess clustering quality, we employed two complementary metrics, the Silhouette Score (65) and the Davies-Bouldin Index (DBI) (66). The Silhouette Score quantifies how well a sample fits within its assigned cluster compared to the nearest alternative cluster. For a data point $z_{i}$, the score is defined as
$s(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max\{a(z_{i}),\,b(z_{i})\}},$ (5)
$DBI=\frac{1}{K}\sum_{i=1}^{K}\max_{j\neq i}\left(\frac{s_{i}+s_{j}}{d_{ij}}\right),$ (6)
where $s_{i}$ is the average distance between points in cluster $i$ and its centroid, and $d_{ij}$ is the distance between the centroids of clusters $i$ and $j$. Lower DBI values indicate better clustering performance, favoring compact and well-separated clusters. The detailed algorithm settings for GMM and K-Means are provided in the Supplementary Material.
Uncertainty analysis
@@ Conclusions
Associating standard errors to machine learning predictions is not a common practice (or inexistent to the best of our knowledge). To address this issue, in this section, we propose a simple, yet formal, method to associate uncertainties to the predictions of the models. The idea is simple, and follows the standard method of contrasting predicted and true values. Figures 2 and 3 show various predicted values plotted against known properties. As with any other method, we observe the prediction is not perfect, with the data showing some scatter around the known values. When presented in this form, the predicted and true (experimentally measured) values can be linked by a linear fit, with a regression analysis able to provide us with average and error values for the slope and intercept. In brief, predicted and true values form the function
@@
, where $\zeta_{predicted}$ is the predicted property, which in our case can be either the surface tension or the viscosity, $\zeta_{true}$ is the true (measured) value, $m$ is the slope, and $b$ is the value taken by the prediction at a condition where the true value is zero (or the point where the line crosses the y-axis). In a perfect scenario, $m$ would take a value of one, and $b$ the value of zero, i.e. every prediction matching the true value. However, our predictions are not perfect, $m$ is not one and $b$ is not zero. In fact, our true values are also not uncertainty-free as they were obtained experimentally, with error. Propagation of uncertainty theory tells us that the error on the prediction (Eqn. 7) is given by
$\delta\zeta_{predicted}=\sqrt{\left(\zeta_{true}\delta m\right)^{2}+\left(m\delta\zeta_{true}\right)^{2}+\delta b^{2}},$
where $\delta\zeta_{predicted}$ is the error of the predicted value, $\delta\zeta_{true}$ is the experimental error of the true value, $\delta m$ is the error of the slope, and $\delta b$ is the error of the y-axis crossing. Combining Eqns. 7 and Uncertainty analysis we obtain
$\delta\zeta_{predicted}=\sqrt{\left(\frac{\zeta_{predicted}-b}{m}\right)^{2}(\delta m)^{2}+(m\delta\zeta_{true})^{2}+\delta b^{2}}.$
@@
In contrast, the error analysis applied to models 1 and 3, from XGBoot, e.g. Figures 3a and d, produce much more accurate results:

### v2

@@ VI.2 Unsupervised learning results
We employed unsupervised clustering on the feature representations extracted by the convolutional neural network to identify intrinsic groupings in the droplet dynamics, based on the droplet images in our validation dataset. From the behaviour of the Silhouette and DBI metrics across candidate values of $K$ (see Appendix D), we selected $K=5$ for both K-Means and GMM, corresponding to a maximum Silhouette Score. In the case of the DBI, extending $K$ beyond this value did not result in a large improvement, and, as previously noted, a smaller number of clusters is preferable for insight into the underlying physics.
@@
We computed the average decoded shape of all droplets within a cluster to visualize the typical droplet morphology for that cluster. Figure 9c shows these mean shapes for the GMM and K-Means partitions. Both methods readily and consistently identify Cluster 4 as a bulb attached to a very long filament, and Cluster 0 as a wide and rounded bulb coming from a-small-in-comparison nozzle (with no connecting filament). The other three clusters only differentiate in the relative size of the nozzle and its connecting filament. Cluster 1 consists of elongated droplets connected to a small nozzle through a short filament, Cluster 3 is a droplet connected, through a filament, to an equal-in-size nozzle, and Cluster 2 is a droplet connected to a intermediate-in-size filament and nozzle. The close agreement between the average shapes from the two methods confirms that both clustering approaches capture coherent and meaningful droplet shape features.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
@@
Figure 10: Clustering of droplet breakup regimes in the BoBoOhOh and ReReOhOh parameter spaces. Panels (a) and (b) show $K$-means clustering, while panels (c) and (d) present the Gaussian mixture model (GMM) results.
@@
A central question arising from the clustering is whether the unsupervised method effectively identifies, or differentiates, the underlying regimes and the various other factors influencing the dripping and break-up dynamics. Following previous works [25], that have classified the liquid break-up dynamics in terms of scaling arguments, Figure 10a and Figure 10c show the cluster data projected onto the log-log ($Re$, OhOh) plane for GMM and K-Means. Importantly, some clusters are readily physically interpretable. Decreasing $Oh$ indicates weakening viscous effects relative to inertial-capillary stresses, while increasing $Re$ reflects stronger inertia at fixed viscosity. Both methods clearly separate Cluster 4 and Cluster 0. Cluster 4 lies predominantly in the upper-left quadrant (high $Oh$, low $Re$), consistent with viscous-damped pinch-off with limited necking and slow recoil, and visually associated with very long filaments. In contrast, Cluster 0 concentrates in the lower-right region (low $Oh$, high $Re$), characteristic of inertial–capillary breakup producing small, nearly spherical droplets with weak viscous damping. Further analyses can be carried out by reflecting on the distribution of liquid formulations among the clusters, as summarized in the Table 3 and 4 in Appendix E. A clear pattern emerges from these analyses: high viscosity fluids, such as glycerin-rich solutions and high cSt silicone oils, are consistently grouped into Cluster 4 by both GMM and K-Means. These samples correspond to the most elongated droplets, supported by the high ratios $L_{2}/L_{1}$ and the mean shapes shown in Figure 9c. In contrast, low-viscosity and low-surface-tension fluids are more frequently assigned to Cluster 0. These groups correspond to rounder, more stable droplets with minimal tailing.
@@
The lack of a common denominator among data clusters 1, 2, and 3 suggests that a further factor, apart from the flow rate, the nozzle size, and the liquid characteristics, is playing a role in driving the dripping dynamics. Fortunately, dripping is well-known to be governed by a third dimensionless parameter, the Bond number [25]. Figure 10a and 10c show the data and clustering in the BoBoOhOh space, where all five clusters are well separated, indicating distinct regimes governed by the relative importance of gravity ($Bo$) and viscosity ($Oh$). Clusters 0-3 occupy a narrow range of $Oh$ but progressively shift toward larger $Bo$. This increase in $Bo$ indicates that these clusters primarily reflect the gradual strengthening of gravitational forces relative to capillarity, while viscous effects remain comparatively constant. Cluster 0 corresponds to the capillary-dominated regime, where droplets are nearly spherical and detach quasi-statically. The gravitational effect strengthens among Clusters 1-3; even though they have comparable viscosity, the increasing gravitation elongates the filaments. In contrast, Cluster 4 deviates markedly from this $Bo$-driven sequence; it spans to much higher $Oh$ values while maintaining moderate $Bo$, this corresponds to high viscous effects in which surface tension and gravity are both resisted by strong viscous damping.
@@
As described in Section IV.2, the clustering was performed on a feature vector that concatenates droplet-shape descriptors with the experimental settings (nozzle diameter and mean flow speed) and does not explicitly include fluid properties. The data separation in ReReOhOh and BoBoOhOh therefore suggests that these features implicitly encode the inertia–viscosity–gravity balance. Notably, the physical properties of the liquids were not input into the clustering. Thus the separation in Re-Oh and Bo-Oh space indicates that our method successfully uncovered the underlying physics linking the experimental settings to the observed droplets.The strong mapping between physical composition and cluster assignment provides further evidence that the unsupervised clusters are not only mathematically separable but also physically interpretable.
VII Uncertainty analysis
@@
Associating standard errors with machine learning predictions is not a common practice (or nonexistent to the best of our knowledge). To address this issue, in this section, we propose a simple, yet formal, method to associate uncertainties to the predictions of the models. The idea is simple, and follows the standard method of contrasting predicted and true values. Figures 7 and 8 show various predicted values plotted against known properties. As with any other method, we observe the prediction is not perfect, with the data showing some scatter around the known values. When presented in this form, the predicted and true (experimentally measured) values can be linked by a linear fit, with a regression analysis able to provide us with average and error values for the slope and intercept. In brief, predicted and true values form the function
@@
where $\zeta_{predicted}$ is the predicted property, which in our case can be either the surface tension or the viscosity, $\zeta_{true}$ is the true (measured) value, $m$ is the slope, and $b$ is the value taken by the prediction at a condition where the true value is zero (or the point where the line crosses the y-axis). In a perfect scenario, $m$ would take a value of one, and $b$ the value of zero, i.e. every prediction matching the true value. However, our predictions are not perfect, $m$ is not one and $b$ is not zero. In fact, our true values are also not uncertainty-free as they were obtained experimentally, with error. Propagation of uncertainty theory tells us that the error on the prediction (Eq. 1) is given by
$\delta\zeta_{predicted}=\sqrt{\left(\zeta_{true}\delta m\right)^{2}+\left(m\delta\zeta_{true}\right)^{2}+\delta b^{2}},$ (2)
where $\delta\zeta_{predicted}$ is the error of the predicted value, $\delta\zeta_{true}$ is the experimental error of the true value, $\delta m$ is the error of the slope, and $\delta b$ is the error of the y-axis crossing. Combining Eqns. 1 and 2 we obtain
$\delta\zeta_{predicted}=\sqrt{\left(\frac{\zeta_{predicted}-b}{m}\right)^{2}(\delta m)^{2}+(m\delta\zeta_{true})^{2}+\delta b^{2}}.$ (3)
@@
In contrast, the error analysis applied to models 1 and 3, from XGBoost, e.g. Figures 8a and d, produce much more accurate results:
@@
VIII Conclusions
Previous studies have demonstrated that the simultaneous measurement of surface tension and viscosity is possible using specialized techniques [61], but most traditional characterization protocols rely on separate tensiometers and viscometers, which often demand substantial sample volumes [62] and/or are limited by their measurement range [16]. Our approach provides a single-step measurement of both surface tension and viscosity, alleviating the need for multiple instruments [63]. Our results demonstrate a generalizable route to rapid, multi-parameter characterization of liquids from a single pinch-off snapshot, which is suited for integration into automated or in-line monitoring systems. The findings open up new possibilities in inkjet printing, e.g. real-time characterisation of ink viscosity and/or surface tension during jetting [42, 64], or on biomedical diagnostics with rapid small-volume property checks [65], of on other industrial processes with on-the-fly materials characterization [66].
By making our dataset (and code) openly available to the community, we expect that the database and models will be extended and potentially improved by including numerically obtained results and additional experiments under conditions not attainable in our facilities, thereby broadening validation and improving generalization.